# Searching for Accreting Compact Binaries in eROSITA DR1

### v1

@@ 3.4 GALEX Ultraviolet Data
We utilize the Galaxy Evolution Explorer (GALEX) Data Release 7 (DR7) to match ultraviolet photometry for our 22 candidate sources. Through cross-matching the GALEX DR7 catalog with our sample within a 5′′ radius, we obtain NUV magnitudes for seven sources, including J0632 and J0859. Additionally, we query the Multi-mission Archive at STScI (MAST) for GALEX tiles upon our candidates and download the associated Coadded Imaging Products. The Merged Catalog of Sources (AIS__mcat.fits) provides NUV magnitudes for other six sources. We note that some faint sources (NUV 23 in the merged catalog) have fluxes approaching the background noise level. for For five faint candidates, we derive flux upper limits of 5 using the original Imaging files (AIS__int.fits). Figure 4 illustrates the distribution of the NUV-to-X-ray flux ratio versus orbital period for our sample. For compact binaries with comparable accretion rates, neutron star or black hole systems are expected to exhibit harder X-ray emission compared to white dwarf systems. The NUV-to-X-ray flux ratios of our candidates span several orders of magnitude. We suggest that sources with higher flux ratios are more likely to be neutron star or black hole X-ray binaries. With additional classified white dwarf and neutron star/black hole systems, we anticipate a clearer separation in this distribution, enabling more precise classification of candidate sources.
@@
Note. — This table summarizes the properties of 22 candidate compact binaries. Columns include: Gaia source designation; parallax () and its signal-to-noise ratio (); Gaia G-band magnitude () and color (); X-ray flux () and luminosity () from eROSITA; orbital period from ZTF light curves; X-ray spectral index from power-law fitting (if available). A dash (–) indicates insufficient photon counts to perform a reliable fit.
@@ 4 Conclusions and Discussion
We combine Gaia DR3 and ZTF time-domain photometry to search for compact binary candidates in the eRASS1 catalog. We focus on those eRASS1 X-ray sources, which match only one Gaia counterparts within a 10′′ radius. Using the optical-to-X-ray flux ratio and Gaia color, we select the X-ray sources whose X-ray emission dominately produced by accretion. We identify 324 sources exhibiting periodic light curves by searching for periodic signals in ZTF time-domain photometry. Finally, we obtain 22 compact binaries candidates after visual inspection.
Among our candidates, 12 sources are classified as RR stars two well-studied compact binaries J0419 and J1023 are included. J0419 is a pre-extremely low mass (pre-ELM) system, and J1023 is a neutron star X-ray binary.
@@
We collect spectroscopic data from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and Sloan Digital Sky Survey (SDSS) for five sources: J0419, J1023, J0632, J0859, and J0945. J0419 and J1023 have multiple spectra, with features consistent with their established classifications. J0632, J0859, and J0945 exhibit spectral characteristics of CV and are classified as CVs by LAMOST and/or SDSS. The Gaia EDR3 catalog provides photometric classifications for 12 additional sources, including CVs, RR Lyrae stars, and eclipsing binaries, as shown in Figure 2. We fit eROSITA X-ray spectra, with J0937 exhibiting a relatively hard spectral type due to higher photon counts. Using ultraviolet photometry, we find that J0937 has a high ultraviolet-to-X-ray flux ratio, suggesting it is likely a neutron star or black hole X-ray binary.
@@
Reflecting on the selection process, requiring a unique Gaia optical counterpart within 10′′ is a stringent criterion. In the absence of an official eRASS1 optical counterpart catalog, this approach effectively identifies accurate optical counterparts, though it excludes a significant number of X-ray sources. As an initial effort in our series of studies, we plan to adopt more relaxed matching criteria in future searches. When excluding active stars, we focus on the optical-to-X-ray flux ratio rather than absolute flux values, simplifying calculations such as X-ray flux conversions. However, this simplification shifts the empirical relation in Figure 2, potentially excluding one additional source. We retain the original boundary in our results, noting that sources near this boundary warrant further analysis. Nevertheless, the adopted boundary effectively distinguishes active stars from accretion-powered systems.
@@
In selecting ZTF periodic variability, we retain sources with clear light curves regardless of their variability patterns. Accretion disks are luminous and unstable, leading to complex light curves in accreting systems that may resemble those of eclipsing binaries or pulsating variables. However, the prior exclusion of active stars using the optical-to-X-ray flux ratio significantly reduces the likelihood of contamination by eclipsing binaries or pulsating variables. The short-period variability sources we identify are highly promising X-ray binary candidates. We plan to request spectroscopic observation time with large telescopes to classify the compact objects and measure their masses through follow-up spectroscopy.
@@
We initiated this study with Gaia DR3 as the optical catalog, representing a preliminary effort to mine compact binaries from eROSITA data. However, as demonstrated throughout this work, our selection does not rely on Gaia astrometric parameters, as no parallax, proper motion, or RUWE cuts were applied. Importantly, the only requirement on the optical catalog is the availability of broad-band photometry to compute colors and flux ratios. This makes our approach fully adaptable to other wide-field photometric surveys with similar depth and coverage. In conjunction with X-ray fluxes and time-domain photometry, this strategy offers a scalable pathway to uncover quiescent LMXBs and CVs beyond Gaia’s astrometric reach.
@@ acknowledgments
This work was supported by the National Natural Science Foundation of China under grant 12433007. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. ZTF is a public-private partnership, with equal support from the ZTF Partnership and from the U.S. National Science Foundation through the Mid-Scale Innovations Program (MSIP).

### v2

@@ 3.4 GALEX Ultraviolet Data
We utilize the Galaxy Evolution Explorer (GALEX) Data Release 7 (DR7) to match ultraviolet photometry for our 22 candidates. Through cross-matching the GALEX DR7 catalog with our sample within a 5′′ radius, we obtain NUV magnitudes for seven sources, including J0632 and J0859. Additionally, we query the Multi-mission Archive at STScI (MAST) for GALEX tiles upon our candidates and download the associated Coadded Imaging Products. The Merged Catalog of Sources (AIS__mcat.fits) provides NUV magnitudes for other six sources. We note that some faint sources (NUV 23 in the merged catalog) have fluxes approaching the background noise level. For four faint candidates, we derive flux upper limits of 5 using the original Imaging files (AIS__int.fits). Figure 5 illustrates the distribution of the NUV-to-X-ray flux ratio versus orbital period for our candidates. The observed flux ratios span several orders of magnitude. We have similarly cross-matched the CVs and MS-LMXBs appearing in Figure 5 with the GALEX DR7 catalog, and plotted them on the figure using the same markers for direct comparison. As with X-ray hardness ratios, differences in accretion physics between neutron star/black hole and white dwarf systems may also appear in softer-band flux ratios, such as NUV-to-X-ray. In principle, this ratio could serve as a boundary between the two populations. However, due to the limited depth of GALEX and the limited size of known sample, a comprehensive analysis has yet to be carried out. With more classified white dwarf and neutron star/black hole systems, we anticipate a clearer separation in this parameter space. This will facilitate more effective classification of candidates in our future work.
@@
Note. — This table summarizes the properties of 22 accreting compact object binary candidates. Columns include: Gaia source designation; parallax () and parallax over error (); Gaia G-band magnitude () and color (); X-ray flux () from eROSITA; X-ray luminosity () estimated from and ; photometric period () from ZTF light curves; X-ray photon index () from power-law fitting to the eROSITA spectrum. A dash (–) indicates insufficient photon counts to fit.
@@ 4 Conclusions and Discussion
Combining Gaia DR3 and ZTF time-domain photometry, we search for accreting compact object binaries in the eRASS1 catalog. We focus on those eRASS1 X-ray sources, which match only one Gaia counterparts within a 10′′ radius. Using the X-ray-to-optical flux ratio and Gaia color, we select the X-ray sources whose X-ray emission dominately produced by accretion. We identify 324 sources exhibiting periodic light curves by searching for periodic signals in ZTF time-domain photometry. Finally, we obtain 22 accreting compact object binary candidates after visual inspection. Among our candidates, two well-studied accreting compact object binaries are included (J0419 and J1023). J0419 is an extremely low mass pre-whtie dwarf binary system, and J1023 is a neutron star X-ray binary.
@@
We collect spectroscopic data from LAMOST and SDSS for five sources: J0419, J0632, J0859, J0945 and J1023. J0419 and J1023 have multiple spectra, with features consistent with their established classifications. J0632, J0859, and J0945 exhibit spectral characteristics of CV and are classified as CVs by LAMOST and/or SDSS. The Gaia EDR3 catalog provides photometric classifications for 12 additional sources, including CVs, RR Lyrae stars, and eclipsing binaries, as shown in Figure 2. We fit eROSITA X-ray spectra using absorbed power-law model, with J0937 exhibiting a relatively hard spectrum, as shown in Figure 1.
@@
Reflecting on the selection process, requiring a unique Gaia optical counterpart within 10′′ is a stringent criterion. In the absence of an official eRASS1 optical counterpart catalog, this approach effectively identifies accurate optical counterparts, though it excludes a significant number of X-ray sources. As an initial effort in our series of studies, we plan to adopt more relaxed matching criteria in future searches. When excluding active stars, we focus on the X-ray-to-optical flux ratio rather than absolute flux values, simplifying calculations such as X-ray flux conversions. However, this simplification shifts the empirical relation in Figure 2, potentially excluding one additional source. We retain the original boundary in our results, noting that sources near this boundary warrant further analysis. Nevertheless, the adopted boundary effectively distinguishes active stars from accretion-powered systems. Overall, we adopted a relatively stringent selection strategy throughout the filtering process, which excluded the majority of sources and yielded only 22 candidates out of over a million initial targets. With the future release of time-domain observations from WFST and future eROSITA data releases, our approach is expected to become increasingly effective in the search for accreting compact object binaries.
@@
In selecting ZTF periodic variability, we retain sources with clear light curves regardless of their variability patterns. Accretion disks are luminous and highly variable, producing complex light curves with features resembling eclipses or pulsations. However, the prior exclusion of active stars using the X-ray-to-optical flux ratio significantly reduces the likelihood of contamination by eclipsing binaries or pulsating variables. The short-period variability sources we identify are highly promising X-ray binary candidates. We obtained spectroscopic data of J0937 using the CFHT, aiming to confirm its accreting nature; the data are currently being reduced. Spectroscopic confirmation of additional reported candidates will be carried out progressively in future observations.
@@
We initiated this study with Gaia DR3 as the optical catalog, representing a preliminary effort to mine accreting compact object binaries from eROSITA data. However, as demonstrated throughout this work, our selection does not rely on Gaia astrometric parameters, as no parallax, proper motion, or RUWE cuts were applied. Importantly, the only requirement on the optical catalog is the availability of broad-band photometry to compute colors and flux ratios. This makes our approach fully adaptable to other other wide-field photometric surveys with enough depth and coverage such as the Rubin Observatory LSST. In conjunction with X-ray-to-optical flux ratio and time-domain photometry, this strategy offers a scalable pathway to uncover quiescent LMXBs and CVs beyond Gaia’s astrometric reach.
@@ acknowledgments
We thank the anonymous referee for the constructive suggestions to improve the paper. This work was supported by the National Natural Science Foundation of China under grants 12433007, 12221003 and 12393814. J.Z.L was supported by the Tianshan Talent Training Program through the grant 2023TSYCCX0101. We acknowledge the science research grants from the China Manned Space Project with No. CMS-CSST-2025-A13. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. ZTF is a public-private partnership, with equal support from the ZTF Partnership and from the U.S. National Science Foundation through the Mid-Scale Innovations Program (MSIP).